# ReGAL: Refactoring Programs to Discover Generalizable Abstractions

### v1

@@ 3.3 Testing
ReGAL-augmented Agent. Our agent has access to the training data and code bank, as well as examples of refactored programs in the demo bank. The ICL budget (10 examples for all experiments) is split between primitive training examples and refactored ones. [^3] We found it necessary to keep some primitive programs as ICL examples, as not all test queries can be handled by helper functions. We treat this ratio of primitive and refactored programs as a hyperparameter. In addition to these demonstrations, the augmented agent retrieves up to 20 relevant helper functions from the code bank, where relevance is measured by the similarity between the query and the function name and description. These helper functions are concatenated into the prompt. The final input is a prompt containing the instructions, the retrieved helper functions, the mixed ICL examples, and the test query. To encourage the model to use helper functions, we include a ReAct-style prompt (Yao et al., 2023) that first asks the model to think about which functions might be relevant based on the query and then generate the code. [^4] Without these additional “thought” statements, we found the augmented agent rarely uses any helper functions.
@@ 4.1 Datasets
We explore three datasets: LOGO, Date understanding, and TextCraft. A common thread through these datasets is that they contain heterogenous problems requiring multiple helper functions as opposed to problems like sorting, which are challenging for LLMs but can be solved with a single function (Dziri et al., 2023). Statistics for the datasets are given in Table 4. Note that the number of primitives reflects functions not built into Python – all models also have access to all Python functions. See Sec. A.5 for further details about each dataset and its primitive operations.
LOGO.   LOGO is based on the Logo Turtle graphics domain-specific language (Abelson & DiSessa, 1986), with which basic graphics can be drawn by controlling a pen (the “turtle”) that draws as it moves through space, using commands like $\tt{forward(dist)}$ and $\tt{left(theta)}$. The data we use is based on the Ellis et al. (2021)’s LOGO dataset, re-annotated by Wong et al. (2021). For easier prediction by LLMs, we parse the data into abstract syntax trees and write a set of rules for translating these into Python; we release this rewritten data. We use the “small” train/test splits (200/111) from Wong et al. (2021) and take $100$ dev examples from the “large” train set.
@@
Table 1: Comparison to relevant past work in each domain. ${}^{\dagger}$ denotes the TextCraft dataset.
@@
LOGO Date TextCraft
Agent Prim. ReGAL Prim. ReGAL Prim. ReGAL
CodeLlama-7B $34.5_{\pm\text{ 1.3}}$ $34.5_{\pm\text{ 1.6}}$ $52.4_{\pm\text{ 0.7}}$ $55.2_{\pm\text{ 1.4}}$ $12.8_{\pm\text{ 1.3}}$ $16.7_{\pm\text{ 1.3}}$
CodeLlama-13B $45.6_{\pm\text{ 0.3}}$ $\mathbf{57.1_{\pm\text{ 0.6}}}$ $42.8_{\pm\text{ 2.0}}$ $68.9_{\pm\text{ 1.6}}$ $18.8_{\pm\text{ 0.7}}$ $26.9_{\pm\text{ 2.2}}$
CodeLlama-34B $50.2_{\pm\text{ 0.8}}$ $50.8_{\pm\text{ 0.6}}$ $47.2_{\pm\text{ 1.5}}$ $68.5_{\pm\text{ 2.1}}$ $22.2_{\pm\text{ 0.7}}$ $\mathbf{30.8_{\pm\text{ 1.3}}}$
Lemur-70B $44.1_{\pm\text{ 1.4}}$ $56.8_{\pm\text{ 0.9}}$ $68.2_{\pm\text{ 0.4}}$ $70.5_{\pm\text{ 0.6}}$ $15.7_{\pm\text{ 1.7}}$ $23.5_{\pm\text{ 2.1}}$
GPT-3.5-turbo $36.9_{\pm\text{ 1.6}}$ $49.3_{\pm\text{ 1.1}}$ $88.9_{\pm\text{ 0.3}}$ $\mathbf{90.2_{\pm\text{ 0.5}}}$ $15.4_{\pm\text{ 1.3}}$ $18.4_{\pm\text{ 2.0}}$
Table 1: Comparison to relevant past work in each domain. ${}^{\dagger}$ denotes the TextCraft dataset.
Table 2: Accuracy of baseline agents predicting primitive programs (Prim.) and those augmented with ReGAL helper functions (3 random seeds). Across domains and models, ReGAL improves over a strong baseline programming agent with access to the same number of ICL examples.
4.2 Baselines
Baselines from Prior Work.   We compare ReGAL against relevant external baselines from past work. However, note that multiple methodological differences in our work, like the use of ICL examples and the format of the output programming language, give our agent an inherent advantage over these baselines. Thus, we refer to these numbers primarily to highlight the strength of our baseline agent. For LOGO, we use the “offline synthesis” numbers reported by Grand et al. (2023), which resemble our train/test setting; however, we note that Grand et al. (2023) predict programs in their original Lisp format and use a different agent model. For the Date dataset, we run Lyu et al. (2023)’s Faithful-CoT method on our custom test split using $\tt{gpt\text{-}3.5\text{-}turbo}$. While the output format and models used are the same, both our baseline and ReGAL use retrieved examples for in-context learning, while Lyu et al. (2023) do not. Furthermore, our ICL examples are based on filtered programs generated by Lyu et al. (2023), leading to better performance even from our baseline agent. Finally, for TextCraft we re-run Prasad et al. (2023)’s baseline – based on ReAct (Yao et al., 2023) – on the depth-2 test set of TextCraft. Here, we use $\tt{gpt\text{-}3.5\text{-}turbo\text{-}instruct}$, as Prasad et al. (2023) found it to outperform $\tt{gpt\text{-}3.5\text{-}turbo}$.
Baseline Programming Agent.   For a more direct comparison, we implement a baseline agent that has access to all the same data as ReGAL but does not use abstractions, thus directly testing the role of ReGAL abstractions in performance. Our baseline agent retrieves primitive programs from the training data; note that this is exactly the same dataset used for refactoring, i.e. the baseline LOGO agent retrieves demonstrations from the LOGO training examples. The input to the baseline agent is a prompt with the same overall instructions as the ReGAL agent (including a description of the primitives), the ICL examples, and the test query; the output is a program for the test query. We use a fixed budget of 10 ICL examples so that the baseline agent sees exactly as many demonstrations as the ReGAL agent.
5 Results
Comparison to External Baselines. Table 2 shows a comparison of the baseline and ReGAL agents the external baselines from prior work. We first note that ReGAL outperforms the baselines in all cases. Furthermore, because of the methodological differences detailed in Sec. 4, our baseline “primitive-only” agent – equipped with ICL examples and using a code-specific LLM – also outperforms past baselines on LOGO and Date. On TextCraft, the ReAct baseline from Prasad et al. (2023) has an advantage in that it receives environmental feedback, while our baseline does not. Nevertheless, even without feedback ReGAL outperforms ReAct. Thus, we compare primarily against our baseline agent, as this provides a direct measure of the impact of abstractions (rather than the other changes made).
@@ 5 Results
ReGAL outperforms the baseline agent. Table 2 shows ReGAL’s performance compared to the baseline agent using primitive programs (described in Sec. 4.2). Overall, for each model type, ReGAL generally outperforms the baseline by a wide margin; for example, ReGAL provides CodeLlama-13B a $11.5\%$ boost on LOGO, allowing it to outperform much larger models. Across datasets, CodeLlama-13B generally benefits most from ReGAL abstractions. Table 2 also shows that large models also benefit from ReGAL, with large gains for Lemur-70B and GPT-3.5 on LOGO and TextCraft. Finally, the largest models are not necessarily the best: on LOGO and TextCraft, GPT-3.5 is outperformed by open-source models, especially after augmentation, e.g. CodeLlama-13B with ReGAL abstractions is able to outperform GPT-3.5 without abstractions by $19.2\%$ (it also outperforms GPT-3.5 with abstractions). Thus, by running ReGAL’s training process on only $\sim\!200$ examples, we can bring a much smaller open-source model’s accuracy far beyond that of a (likely) much larger system.
@@
Ablations. In Sec. 3.2 we describe ReGAL’s multiple components; here we determine the utility of each by removing each one in isolation. Table 3 shows the results of these ablations. We use the CodeLlama-13B agent due to the size of ReGAL’s impact on it across tasks. We average the performance across 3 seeds. Table 3 shows that each component contributes to performance, with drops when any is removed. Across datasets, the largest performance decreases come with the removal of retrials and with removal of the curriculum. Retrials can not only increase the number of useful helper functions but can also help increase the number of examples in the Demo Bank. Replacing the curriculum with a random ordering also severely hurts performance, e.g. leading to an $18.7\%$ drop on LOGO.
@@
ReGAL learns general, reusable functions. In Sec. 1, we stressed the importance of reusability. Specifically, generating programs without shared abstractions means that the model has to re-generate subprograms that could be reused across multiple test instances. We argue that ReGAL improves over this paradigm by learning shared abstractions. The results in Table 2 indicate that ReGAL offers large improvements over a baseline agent that lacks abstractions. Here, we verify that the abstractions learned are reusable, i.e. shared. Fig. 3 shows the number of times the top-5 most common ReGAL functions are called in test programs produced by the CodeLlama-13B agent. Across all datasets, we see that the helper functions learned by ReGAL are commonly reused, with the most relative reuse in TextCraft. Sec. B.1 shows examples of these common functions.
@@ 6.1 What kinds of programs are discovered?
To further examine what kinds of helper functions are discovered, we examine the most frequent helper functions for each domain from Fig. 3, summarizing the results below. Refer to Sec. B.1 for function code. We find that distinct trends emerge across domains.
For LOGO, ReGAL discovers functions that encapsulate different types of shapes. This is expected, as the LOGO data was generated with these functionalities in mind, i.e. the larger shapes are composed of objects like semicircles, pentagons, and circles. For Date, ReGAL tends to encapsulate single operations, prioritizing interpretability with names like $\tt{get\_date\_one\_year\_ago\text{()}}$). While seemingly less complex than LOGO’s functions, this approach aligns with the importance of function naming in synthesis procedures, as highlighted by Grand et al. (2023). In TextCraft, the functions uncovered by ReGAL are more complex and reflect the dynamics of the game. Specifically, the functions include conditional statements for checking ingredients, reflecting the fact that in TextCraft, having the correct crafting ingredients is a prerequisite for crafting an object (see $\tt{craft\_and\_get\_ingredient\text{()}}$) in Fig. 2 and Table 10, which is taken from the learned code bank $C$).
Figure 3: Function usage for CodeLlama-13B for the top-5 most common helpers. Functions are reused across programs.
6.2 Error Analysis
@@
Thus, the most common error is a failure to predict primitive operations; here, the ReGAL agent is at a disadvantage w.r.t. the baseline agent, since both have the same ICL budget. The baseline agent sees 10 ICL examples with only primitive code, while the ReGAL agent sees 5 primitive examples and 5 Demo Bank examples.
@@
6.3 Sample Efficiency
@@
7 Discussion and Conclusion

### v2

@@ 3.3 Testing
ReGAL-augmented Agent. Our agent has access to the training data and code bank, as well as examples of refactored programs in the demo bank. The ICL budget (10 examples for all experiments) is split between primitive training examples and refactored ones. [^3] We found it necessary to keep some primitive programs as ICL examples, as not all test queries can be handled by helper functions alone. We treat the ratio of primitive to refactored programs in the ICL example as a hyperparameter (all values listed in Table 9). In addition to these demonstrations, the augmented agent retrieves up to 20 relevant helper functions from the code bank, where relevance is measured by the similarity between the query and the function name and description. These helper functions are concatenated into the prompt. The final input is a prompt containing the instructions, the retrieved helper functions, the mixed ICL examples, and the test query. To encourage the model to use helper functions, we include a ReAct-style prompt (Yao et al., 2023) that first asks the model to think about which functions might be relevant based on the query and then generate the code. [^4] Without these additional “thought” statements, we found the augmented agent rarely uses any helper functions.
@@ 4.1 Datasets
We explore five datasets: LOGO, Date understanding, TextCraft, MATH, and TabMWP. A common thread through these datasets is that they contain heterogenous problems requiring multiple helper functions as opposed to problems like sorting, which are challenging for LLMs but can be solved with a single function (Dziri et al., 2023). Statistics for the datasets are given in Table 5. See Sec. A.5 for details about each dataset and its primitive operations.
LOGO.   LOGO is based on the Logo Turtle graphics domain-specific language (Abelson & DiSessa, 1986), with which basic graphics can be drawn by controlling a pen (the “turtle”) that draws as it moves through space, using commands like $\tt{forward(dist)}$ and $\tt{left(theta)}$. The data we use is based on Ellis et al. (2021)’s LOGO dataset, re-annotated by Wong et al. (2021). For easier prediction by LLMs, we parse the data into abstract syntax trees and write a set of rules for translating these into Python; we release this rewritten data. We use the “small” train/test splits (200/111) from Wong et al. (2021) and take $100$ dev examples from the “large” train set.
@@
Table 1: Accuracy of baseline agents predicting primitive programs (Prim.) and those augmented with ReGAL helper functions (3 random seeds). Across domains and models, ReGAL improves over a strong baseline agent with access to the same number of ICL examples. Math domains with no clear domain language marked with .
LOGO Date TextCraft MATH (Alg.) TabMWP
Agent Prim. ReGAL Prim. ReGAL Prim. ReGAL Prim. ReGAL Prim. ReGAL
CodeLlama-7B $34.5_{\pm\text{ 1.3}}$ $34.5_{\pm\text{ 1.6}}$ $52.4_{\pm\text{ 0.7}}$ $55.2_{\pm\text{ 1.4}}$ $12.8_{\pm\text{ 1.3}}$ $16.7_{\pm\text{ 1.3}}$ $6.8_{\pm\text{ 0.8}}$ $17.6_{\pm\text{ 0.8}}$ $16.7_{\pm\text{ 1.6}}$ $27.7_{\pm\text{ 0.6}}$
CodeLlama-13B $45.6_{\pm\text{ 0.3}}$ $\mathbf{57.1_{\pm\text{ 0.6}}}$ $42.8_{\pm\text{ 2.0}}$ $68.9_{\pm\text{ 1.6}}$ $18.8_{\pm\text{ 0.7}}$ $26.9_{\pm\text{ 2.2}}$ $14.0_{\pm\text{ 0.9}}$ $17.6_{\pm\text{ 0.8}}$ $29.1_{\pm\text{ 0.6}}$ $27.9_{\pm\text{ 0.4}}$
CodeLlama-34B $50.2_{\pm\text{ 0.8}}$ $50.8_{\pm\text{ 0.6}}$ $47.2_{\pm\text{ 1.5}}$ $68.5_{\pm\text{ 2.1}}$ $22.2_{\pm\text{ 0.7}}$ $\mathbf{30.8_{\pm\text{ 1.3}}}$ $14.9_{\pm\text{ 1.1}}$ $22.5_{\pm\text{ 1.6}}$ $19.4_{\pm\text{ 1.2}}$ $27.0_{\pm\text{ 0.8}}$
Lemur-70B $44.1_{\pm\text{ 1.4}}$ $56.8_{\pm\text{ 0.9}}$ $68.2_{\pm\text{ 0.4}}$ $70.5_{\pm\text{ 0.6}}$ $15.7_{\pm\text{ 1.7}}$ $23.5_{\pm\text{ 2.1}}$ $14.0_{\pm\text{ 1.2}}$ $13.1_{\pm\text{ 1.2}}$ $27.5_{\pm\text{ 0.9}}$ $25.7_{\pm\text{ 0.8}}$
GPT-3.5-turbo $36.9_{\pm\text{ 1.6}}$ $49.3_{\pm\text{ 1.1}}$ $88.9_{\pm\text{ 0.3}}$ $\mathbf{90.2_{\pm\text{ 0.5}}}$ $15.4_{\pm\text{ 1.3}}$ $18.4_{\pm\text{ 2.0}}$ $43.2_{\pm\text{ 1.6}}$ $\mathbf{55.4}_{\pm\text{ 2.8}}$ $87.4_{\pm\text{ 0.9}}$ $\mathbf{89.2}_{\pm\text{ 0.8}}$
MATH.   To test ReGAL’s ability to discover functions in domains that do not have a pre-defined domain language, we additionally examine the MATH dataset (Hendrycks et al., 2021) consisting of challenging math word problems. As with Date, we generate silver training programs; in this case, we use GPT-4 with a Program-of-Thoughts prompt (Chen et al., 2022). To have a good chance of generating sufficient numbers of training programs, we use examples from hardness levels 1 and 2 (MATH contains 5 levels of difficulty). We also focus on the Algebra subset of MATH, which is the largest. This gives us a train/dev/test split of 194/61/74. While only correct programs are used for training, correct and incorrect examples are used in dev and test. Note that unlike the other datasets, MATH does not have an inherent domain language associated with it and does not have any primitives beyond the ones already contained in Python.
TabMWP.   We further extend our general experiments on MATH by testing on TabMWP (Lu et al., 2022), a tabular resoning dataset consisting of math word problems about tabular data. Here, we also use silver programs for training and focus on levels 1-4 (TabMWP has 8 difficulty levels), following a similar procedure as for MATH. This gives us a train/dev/test split of 194/60/74.
4.2 Baselines
Baselines from Prior Work.   We compare ReGAL against relevant external baselines from past work. However, note that multiple methodological differences in our work, like the use of ICL examples and the format of the output programming language, give our agent an inherent advantage over these baselines. Thus, we refer to these numbers primarily to highlight the strength of our baseline agent. For LOGO, we use the “offline synthesis” numbers reported by Grand et al. (2023), which resemble our train/test setting; however, we note that Grand et al. (2023) predict programs in their original Lisp format and use a different agent model. For the Date dataset, we run Lyu et al. (2023)’s Faithful-CoT method on our custom test split using $\tt{gpt\text{-}3.5\text{-}turbo}$. While the output format and models used are the same, both our baseline and ReGAL use retrieved examples for in-context learning, while Lyu et al. (2023) do not. Furthermore, our ICL examples are based on programs generated by Lyu et al. (2023) after filtering for correctness, leading to better performance even from our baseline agent. Finally, for TextCraft we re-run Prasad et al. (2023)’s baseline – based on ReAct (Yao et al., 2023) – on the depth-2 test set of TextCraft. Here, we use $\tt{gpt\text{-}3.5\text{-}turbo\text{-}instruct}$, as Prasad et al. (2023) found it to outperform $\tt{gpt\text{-}3.5\text{-}turbo}$.
Baseline Programming Agent.   For a more direct comparison, we implement a baseline agent that has access to all the same data as ReGAL but does not use abstractions, thus directly testing the role of ReGAL abstractions in performance. Our baseline agent retrieves primitive programs from the training data; note that this is exactly the same dataset used for refactoring, i.e., the baseline LOGO agent retrieves demonstrations from the LOGO training examples. The input to the baseline agent is a prompt with the same overall instructions as the ReGAL agent (including a description of the primitives), the ICL examples, and the test query; the output is a program for the test query. We use a fixed budget of 10 ICL examples so that the baseline agent sees exactly as many demonstrations as the ReGAL agent.
5 Results
Comparison to External Baselines. Table 2 shows a comparison of the baseline and ReGAL agents the external baselines from prior work. Note that we do not include MATH and TabMWP here as we use a subset of levels for these datasets. We first note that ReGAL outperforms the baselines in all cases. Furthermore, because of the methodological differences detailed in Sec. 4, our baseline “primitive-only” agent – equipped with ICL examples and using a code-specific LLM – also outperforms past baselines on LOGO and Date. On TextCraft, the ReAct baseline from Prasad et al. (2023) has an advantage in that it receives environmental feedback, while our baseline does not. Nevertheless, even without feedback ReGAL outperforms ReAct. Thus, we compare primarily against our baseline agent, as this provides a direct measure of the impact of abstractions (rather than the other changes made).
@@ 5 Results
Table 2: Comparison to relevant past work in each non-math domain. TC denotes the TextCraft dataset.
@@
ReGAL outperforms the baseline agent in non-math domains. Table 1 shows ReGAL’s performance compared to the baseline agent using primitive programs (described in Sec. 4.2). Overall, for each model type, ReGAL generally outperforms the baseline by a wide margin; for example, ReGAL provides CodeLlama-13B a $11.5\%$ boost on LOGO, allowing it to outperform much larger models. Across datasets, CodeLlama-13B generally benefits most from ReGAL abstractions. Table 1 also shows that large models also benefit from ReGAL, with large gains for Lemur-70B and GPT-3.5 on LOGO and TextCraft. Finally, the largest models are not necessarily the best: on LOGO and TextCraft, GPT-3.5 is outperformed by open-source models, especially after augmentation, e.g., CodeLlama-13B with ReGAL abstractions is able to outperform GPT-3.5 without abstractions by $19.2\%$ (it also outperforms GPT-3.5 with abstractions). Thus, by running ReGAL’s training process on only $\sim\!200$ examples, we can bring a much smaller open-source model’s accuracy far beyond that of a (likely) much larger system.
ReGAL improves over the baseline on mathematical domains. Our math domains (MATH and TabMWP) differ from the others in that there is no clear domain language that ReGAL should discover. Unlike a domain like LOGO (where it is clear which functions a human programmer would write), there is less clarity on which helper functions should be discovered. Nevertheless, in this challenging setting ReGAL generally improves over the baseline agent across different models, with major gains for GPT-3.5 on MATH ($11.8\%$) and for CodeLlama-7B on TabMWP ($11\%$), where the 7B ReGAL performance is comparable to the 13B performance. Overall, Lemur struggles with these problems, producing a large number of empty outputs both for the baseline and ReGAL agents; when taking the standard error into account, baseline and ReGAL performance is roughly comparable. These results highlight ReGAL’s flexibility and ability to generalize to domains without clear-cut domain languages.
@@
Ablations. In Sec. 3.2 we describe ReGAL’s multiple components; here we determine the utility of each by removing each one in isolation for non-math datasets. Table 3 shows the results of these ablations. We use the CodeLlama-13B agent due to the size of ReGAL’s impact on it across tasks. We average the performance across 3 seeds. Table 3 shows that each component contributes to performance, with drops when any is removed. Across datasets, the largest performance decreases come with the removal of retrials and with removal of the curriculum. Retrials can not only increase the number of useful helper functions but can also help increase the number of examples in the Demo Bank. Replacing the curriculum with a random ordering also severely hurts performance, e.g., leading to an $18.7\%$ drop on LOGO.
@@
ReGAL learns general, reusable functions. In Sec. 1, we stressed the importance of reusability. Specifically, generating programs without shared abstractions means that the model has to re-generate subprograms that could be reused across multiple test instances. We argue that ReGAL improves over this paradigm by learning shared abstractions. The results in Table 1 indicate that ReGAL offers large improvements over a baseline agent that lacks abstractions. Here, we verify that the abstractions learned are reusable, i.e., shared. Fig. 3 shows the number of times the top-5 most common ReGAL functions are called in test programs produced by the CodeLlama-13B agent. Across all datasets, we see that the helper functions learned by ReGAL are commonly reused, with the most relative reuse in TextCraft. Sec. B.1 shows examples of these common functions.
@@ 6.1 What kinds of programs are discovered?
To further examine what kinds of helper functions are discovered, we examine the most frequent helper functions for each domain from Fig. 3, summarizing the results below. Refer to Sec. B.1 for the implementation of these functions. We find that distinct trends emerge across domains.
For LOGO, ReGAL discovers functions that encapsulate different types of shapes. This is expected, as the LOGO data was generated with these functionalities in mind, i.e., the larger shapes are composed of objects like semicircles, pentagons, and circles. For Date, ReGAL tends to encapsulate single operations, prioritizing interpretablity in function names like $\tt{get\_date\_one\_year\_ago\text{()}}$). While seemingly less complex than LOGO’s functions, this approach aligns with the importance of function naming in synthesis procedures, as highlighted by Grand et al. (2023). In TextCraft, the functions uncovered by ReGAL are more complex and reflect the dynamics of the game. Specifically, the functions include conditional statements for checking ingredients, reflecting the fact that in TextCraft, having the correct crafting ingredients is a prerequisite for crafting an object (see $\tt{craft\_and\_get\_ingredient\text{()}}$) in Fig. 2 and Fig. 7, which is taken from the learned code bank $C$).
Figure 3: Function usage by CodeLlama-13B for the top-5 most common helpers illustrating reusability across examples. The x-axis denotes the number of times a functions is used in the test set.
6.2 What kinds of errors do agents make?
@@
Thus, the most common error is a failure to predict primitives; here, the ReGAL agent is at a disadvantage w.r.t. the baseline, as they share the same ICL budget. The baseline agent sees 10 examples with only primitive code, while the ReGAL agent sees 5 primitive and 5 Demo Bank examples.
@@
6.3 How many training examples does ReGAL need?
@@
6.4 How can ReGAL adapt to distribution shifts?
To assess the usefulness of abstractions learned by ReGAL under a distribution shift at inference-time, we use the TextCraft dataset. We redistribute the instances into three splits: train (seen), dev (unseen), test (unseen) consisting of 100/20/76 instances, such that the dev and test sets contain objects never encountered during training. For example, if the test set requires crafting “yellow jungle fence”, the train set does not contain any instance with a “fence” as a target object. Furthermore, we study how to adapt the learned library with unseen examples from the dev set using ReGAL in two ways: (i) only pruning the learned library to retain generalizable functions (Stage 3(b) of ReGAL); and (ii) an additional iteration of ReGAL training.
Table 4: Performance of baseline and ReGAL agent under distribution shift on TextCraft with unseen objects in the test set.
Method Adaptation Setting Acc.
Prim. Programs -- 7.8
ReGAL -- 11.4
ReGAL $\tt{pruneCodeBank}$ on dev (unseen) 14.9
ReGAL All ReGAL stages on dev (unseen) 17.2
The results are presented in Table 4. First, we observe that the performance of both baseline and ReGAL agents degrades under distribution shift because the in-context examples may not be adequately relevant to the test instance. However, ReGAL outperforms the baseline agent with primitive programs by $3.6\%$. Furthermore, simply pruning an existing code bank increases performance of ReGAL agent by an additional $3.5\%$. Note that this still uses a learned codebank from prior training and without proposing any new abstractions. Finally, learning new abstractions and editing existing ones based on unseen dev data is the most effective yielding an additional $2.3\%$ over only performing $\tt{prunceCodeBank}$ on the unseen data for a total of $9.4\%$ improvement over the baseline.
7 Discussion